SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Kim Hojeong, Rahmat Rahmat, Mark Tibbetts.

Operations

  • A couple of incidents of both minor (just start new run) and major (power cycle TSF crates) trigger damage. Mike P. suggested we swap the ROM boards, as they've seen this problem before as a symptom of single board processor issues. Matt Weaver will do this when he has the chance. This may not fix the problem but it would at least rule out ROM problems.

EMT

  • Occasional cable errors show up in QA monitoring plots. Mark will investigate what type of error it is and whether it is causing problems for data-taking.
  • Board is back from Lupe and has been fixed. It's in IR2 and will be tested when the opportunity arises.
  • Efficiency study: will be summarised in collab meeting talk and BaBar note.
    Further studies - scatter plot of ADC counts vs EMC cluster energy for data and MC to check conversion factor.
  • Should check when/if last EMC source calibration constants were uploaded.
  • Should check if EMT is calibrated when EMC standalone calibration is performed.

Data Quality and Monitoring

  • No problems in data so far.
  • Still no ideas about what causes removal of connecting lines on strip charts. Also still need to adjust scale.

Trigger Configuration

  • Hojeong: see summary page for details of progress on statistics study and also filter study.
  • Rainer: see folder for details of status of trigger emulation study. .tcf and *diff* files summerise the different stages of tightened trigger configurations, for those of us who've forgotten over the previous months. .ps files show effect of the different configurations.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.